# Multi-Conditional Ranking with Large Language Models

### v1

@@ 4.3 Accuracy of Decomposition
Now that we have seen how EXSIR improve LLMs’ performance, one remaining question is that how the accuracy of the decomposition step influences the overall performance of the model. We detail the accuracy of LLMs in extracting and sorting conditions (decompostion step) in Table 3. The result indicates that GPT-4 consistently outperforms in these tasks, whereas Mistral’s accuracy decreases when transitioning to paragraph-level, correlating with its EXSIR-augmented ranking performance in such scenarios.
@@
Figure 4: Evaluating the impact of EXSIR against zero-shot CoT prompting for token-level items. We additionally utilize ColBERT as a representative of encoder-based rankers.
@@ 4.4 Zero-shot CoT vs Decomposed Reasoning
So far, we have observed how EXSIR enhances LLM performance and the correlation between the accuracy of the decomposition step and their overall effectiveness. However, one might question if the multi-step decomposition reasoning is essential. Could a similar performance level be achieved by integrating the decomposition steps into a single prompt, akin to the zero-shot Chain-of-Thought (CoT) (Wei et al., 2022) approach? This section narrows the focus to GPT-4, comparing its performance using EXSIR against zero-shot CoT-style prompting. We provide an example for CoT-based prompt in the Appendix.
The accuracy of GPT-4, GPT-4 with CoT and with EXSIR on MCRank, is depicted in Figures 4 and 5. For token-level items, the figures demonstrate that while CoT prompting boosts the base performance of GPT-4, there remains a notable performance disparity between EXICR and CoT, highlighting the value of multi-step reasoning. In contrast, for paragraph-level items, incorporating CoT instructions seems to decrease the base model’s performance, possibly due to the task complexity and the challenge of adhering to the provided CoT instructions for GPT-4.
4.5 Encoder-Based Ranker on MCRank
In this section, we reexamine our preliminary hypothesis that the MCR task poses inherent challenges for smaller encoder-based rankers. To evaluate this, we utilize ColBERT (Khattab & Zaharia, 2020), trained on the MS MARCO passage ranking task (Nguyen et al., 2016), as a representative for encoder-based ranking models. The accuracy results of ColBERT on MCRank are illustrated in Figures 4 and 5. These figures indicate that ColBERT’s performance is significantly inferior compared to GPT-4 (as well as the other LLMs under investigation), underscoring the task’s complexity and the potential limitations of smaller ranking models in tackling such challenges in multi-conditional setting.
Figure 5: Evaluating the impact of EXSIR against zero-shot CoT prompting for paragraph-level items. We additionally utilize ColBERT as a representative of encoder-based rankers.
@@ 5 Related Work
Large language models have achieved significant success in tackling ranking tasks in recent years. However, despite this progress and the broad applicability of ranking in real-world scenarios, these tasks primarily focus on ranking passages in response to a specific query.
Ranking with LLMs In recent years, LLMs have become pivotal in addressing ranking related tasks. Initially focusing on encoder-based rankers (Nogueira et al., 2019; Khattab & Zaharia, 2020), the rapid advancement of autoregressive LLMs has led to the development of methodologies that utilize these models as rankers, achieving unparalleled performance across various benchmarks (Zhuang et al., 2023a; Qin et al., 2023). However, despite these advances, the majority of LLM-based ranking efforts have concentrated on ordering extensive lists of passages based on a query, often overlooking the diverse applications of ranking in real-world scenarios. Our work closely aligns with developments in recommendation systems, such as the conditional methods proposed by Hou et al. (2024), which only considers a limited concept of condition in regard to variety and complexity compared to our notion of multi-conditional ranking.
Decomposed Reasoning with LLMs As LLMs grow stronger, decomposed reasoning has emerged as a fundamental strategy to enhance their capabilities by segmenting complex tasks into smaller, more manageable components. This decomposition can be straightforward, utilizing a single LLM, as seen in approaches like Chain-of-Thought (Wei et al., 2022), Tree-of-Thought (Yao et al., 2024), and Self-Verification (Weng et al., 2022). Alternatively, it can involve more complex interactions among multiple models within a multi-agent systems (Xi et al., 2023; Guo et al., 2024). Prior research has successfully integrated decomposed reasoning into various tasks, including question answering (Dua et al., 2022), retrieval-augmented generation (RAG) (Asai et al., 2023), and mathematical reasoning (Qi et al., 2023).
@@ 6 Conclusion
We present an in-depth exploration of the multi-conditional ranking (MCR) task, a critical yet underexplored aspect of ranking task in real-world applications. Through the introduction of the MCRank benchmark, we have highlighted the challenges LLMs face when tasked with ranking a small set of items under a variety of complex and sometimes conflicting conditions. Our investigation reveals a significant performance drop in LLMs as the number of conditions and items increases. To address this, we proposed a novel decomposed reasoning approach, EXSIR, which significantly boosts LLMs performance on the MCRank, demonstrating up to a 12% improvement in accuracy. Our comprehensive analysis not only sheds light on the performance of LLMs across different condition categories but also underscores the effectiveness of the decomposition step in enhancing model accuracy. Finally, by contrasting our approach with other existing methodologies such as CoT and an encoder-based ranker, we have illustrated the advantages of EXSIR and the intricate nature of the MCR task.
@@ 7 Limitations
Limited Scope of LLMs: Our research focused on three specific LLMs, which, while prominent, do not encompass the full spectrum of models available in the field. This narrow focus may not fully capture the diversity of capabilities present in the broader landscape of large language models.
@@ Appendix A MCRank Details
In this section, we first outline the specifics of the datasets from which we extract items’ label to construct MCRank. Subsequently, we present a comprehensive list of the various conditions included in the MCRank benchmark.
@@ A.1 Benchmark Used To Create MCRank
A.2 Conditions in MCRank
@@
Appendix B Details of Prompts
The prompts utilized for ranking token-level and paragraph-level items are detailed in Prompts B and B, respectively. Additionally, the prompts employed for the extraction and sorting of conditions are outlined in Prompts B and B. Finally, we provide an example of zero-shot CoT-based prompt for token-level items in Prompt B.
Item-1: [item-1]
Item-2: [item-1]
Conditions: [string of conditions]
Conditions: [list of extracted conditions]
@@ Appendix B Details of Prompts
Appendix C Detailed Breakdown of Ranking performance on MCRank
@@
A detailed breakdowns of models performance on MCRank, segmented by the category of conditions and items, are presented in Tables 5, 6, 7, 8, 9, and 10.

### v3

@@ 4.3 Accuracy of Decomposition
Now that we have seen how EXSIR improve LLMs’ performance, one remaining question is that how the accuracy of the decomposition step influences the overall performance. We detail the accuracy of LLMs in extracting and sorting conditions (decompostion step) in Table 2. The result indicates that GPT-4 outperforms other LLMs, whereas Mistral’s accuracy decreases when transitioning to paragraph-level, correlating with its EXSIR-augmented ranking performance in such scenarios.
@@
Figure 4: Evaluating the impact of EXSIR against zero-shot CoT prompting for token-level items. We additionally report SFR and RankGPT performances as representatives of existing rankers.
@@ 4.4 Zero-shot CoT vs Decomposed Reasoning
So far, we have seen how EXSIR enhances LLMs’ performances. However, one might ask: is multi-step decomposition necessary, or could similar results be achieved by combining the steps into a single prompt, like the zero-shot Chain-of-Thought (CoT) approach (Wei et al., 2022)? This section narrows the focus to GPT-4, comparing its performance using EXSIR against zero-shot CoT-style prompting. We provide an example of a CoT-based prompt and discuss the impact of CoT prompting on Llama3.1’s performance, further reinforcing our findings with GPT-4 in the Appendix.
The accuracy of GPT-4, GPT-4 with CoT and with EXSIR on MCRank, is depicted in Figures 4 and 5. For token-level items, the figures demonstrate that while CoT prompting boosts the base performance of GPT-4, there remains a notable performance disparity between EXICR and CoT, highlighting the value of multi-step reasoning. In contrast, for paragraph-level items, incorporating CoT instructions seems to decrease the base model’s performance, possibly due to the task complexity and the challenge of adhering to the provided CoT instructions for GPT-4 (we provide more discussion on the impact of few-shot CoT in Appendix).
Figure 5: Evaluating the impact of EXSIR against zero-shot CoT prompting for paragraph-level items. We additionally report SFR and RankGPT performances as representatives of existing rankers.
4.5 Existing Rankers and o1-mini
In here, we first evaluate the performance of existing rankers, followed by an assessment of the reasoning-focused LLM o1-mini on MCRank.
Existing rankers:
We evaluate the performance of SFR (Meng et al., 2024) and ColBERT (Khattab and Zaharia, 2020), as representative encoder-based ranking models, along with RankGPT (based on GPT-4) (Sun et al., 2023) on MCRank, due to their strong performance on existing benchmarks. The accuracy of SFR and RankGPT on MCRank are illustrated in Figures 4 and 5 (the performance of ColBERT, along with implementation details, is provided in the Appendix). These figures indicate that SFR’s performance is significantly inferior compared to GPT-4 (and most of other investigated LLMs), highlighting the complexity of the task and the limitations of smaller ranking models in a multi-conditional setting. Moreover, since RankGPT is designed to monotonically rank documents based on a given query, it outperforms the GPT-4 baseline when only one condition is present. However, as the number of conditions increases, its performance becomes comparable to our vanilla GPT-4 baseline. Based on these results, integrating RankGPT into EXSIR’s ranking step could improve performance, which we leave for future research.
o1-mini: To further evaluate the complexity of MCRank, we present the performance of the reasoning-focused LLM o1-mini in Figures 4 and 5 (further experimental results on o1-mini’s are provided in the Appendix). Despite its strong performance on other reasoning tasks, o1-mini shows relatively poor performance on the MCRank benchmark. As evident, GPT-4 combined with EXSIR outperforms o1-mini in many cases. However, in the paragraph-level setting, in cases where more items are involved, o1-mini shows an advantage and performs better. Moreover, as we can observe, applying EXSIR on top of o1-mini significantly improves performance, further demonstrating the effectiveness of the EXSIR approach.
@@ 5 Related Work
LLMs have achieved significant success in ranking tasks, but most efforts focus on ranking passages based on a single query, despite the broad real-world applications of ranking.
Ranking with LLMs: In recent years, LLMs have become pivotal in addressing ranking related tasks. Initially focusing on encoder-based rankers (Nogueira et al., 2019; Khattab and Zaharia, 2020), the rapid advancement of autoregressive LLMs has led to the development of methodologies that utilize these models as rankers, achieving unparalleled performance across various benchmarks (Zhuang et al., 2023a; Qin et al., 2023). However, despite these advances, the majority of LLM-based ranking efforts have concentrated on ordering extensive lists of passages based on a query, often overlooking the diverse applications of ranking in real-world scenarios. Our work closely aligns with developments in recommendation systems, such as the conditional methods proposed by Hou et al. (2024), which only considers a limited concept of condition in regard to variety and complexity compared to our notion of multi-conditional ranking.
Decomposed Reasoning with LLMs: As LLMs grow stronger, decomposed reasoning has emerged as a fundamental strategy to enhance their capabilities by segmenting complex tasks into smaller, more manageable components. This decomposition can be straightforward, utilizing a single LLM, as seen in approaches like Chain-of-Thought (Wei et al., 2022), Tree-of-Thought (Yao et al., 2024), and Self-Verification (Weng et al., 2022). Alternatively, it can involve more complex interactions among multiple models within a multi-agent systems (Xi et al., 2023; Guo et al., 2024). Prior research has successfully integrated decomposed reasoning into various tasks, including question answering (Dua et al., 2022), retrieval-augmented generation (RAG) (Asai et al., 2023), and mathematical reasoning (Qi et al., 2023).
@@ 6 Conclusion
We explore the multi-conditional ranking task, an essential but underexplored aspect of ranking in real-world applications. Introducing MCRank benchmark, we have highlighted the challenges LLMs face when ranking a small set of items under a variety of complex and sometimes conflicting conditions. Our investigation reveals a significant performance drop in LLMs as the number of conditions and items increases. To address this, we introduce EXSIR, a novel decomposed reasoning method that improves LLM performance by up to 14.4% in accuracy. We also analyze the performance of LLMs across various condition categories and the effectiveness of the decomposition step in enhancing accuracy. Finally, by contrasting our approach with other existing methodologies such as CoT and existing rankers, we demonstrate its effectiveness and the complexity of the MCR task.
@@ 7 Limitations
Limited Scope of LLMs: Our research focused on five specific LLMs and three existing rankers, which, while prominent, do not encompass the full spectrum of models available in the field. This narrow focus may not fully capture the diversity of capabilities present in the broader landscape of large language models.
@@
EXSIR Efficiency: Our proposed method EXSIR, while effective in enhancing performance on MCR tasks, presents notable challenges in terms of efficiency and cost. As a multi-step ranking process, EXSIR inherently is more time-consuming and costly compared to simpler, single-step methods. This issue becomes more pronounced when deploying EXSIR at scale in real-world applications. Optimizing the efficiency of EXSIR, without compromising its performance benefits, remains an open area for future research.
@@ Appendix A MCRank Details
In this section, we first outline the specifics of the datasets from which we extract items’ label to construct MCRank. Subsequently, we present more details on how we create MCRank and provide a comprehensive list of the various conditions included in the MCRank benchmark.
@@ A.1 Benchmark Used To Create MCRank
A.2 MCRank Creation Details
We provided a step-by-step illustration of MCRank creation in Figure 6. For each scenario in MCRank, we began with 200 samples for each category, utilizing extracted attributes from the original datasets to establish the golden ranking of items. Additional conditions were then applied on top of each category-based condition, necessitating a recalculation of the gold ranking to accommodate these augmented conditions. Throughout this process, we removed some samples where the addition of extra conditions resulted in non-unique golden rankings. Consequently, the average number of samples per category in MCRank ranges from 159 to 200. The statistics of MCRank are presented in Table 3.
1 Condition 2 Conditions 3 Conditions
T-level 916.7 860.0 797.7
P-level 1000 1000 1000
Table 3: Average number of samples in MCRank benchmark per number of conditions for paragraph- (P-level) and token-level (T-level) items.
Figure 6: Step-by-step illustration of MCRank creation.
A.3 Conditions in MCRank
@@
Appendix B Real-World Example Scenarios for MCR
In addition to the provided scenario for the use of MCR in recommendation systems, multi-conditional ranking can have many use-cases in other applications such as:
Education:
Consider a teacher who needs to select questions from a pre-existing list. Without needing to read or solve each question, the teacher might want to rank them based on conditions such as: (1) Questions from a specific topic (Topic X) should appear higher in the list, with low priority; (2) The questions should be sorted by difficulty, with medium priority; and (3) The most difficult questions should be placed at the end, with high priority, as the teacher prefers to reserve them for advanced students.
HR:
A recruiter is tasked with selecting a candidate from a list of already short-listed resumes. Without delving deeply into every resume, and spending a significant amount of time, the recruiter wants to rank the candidates according to conditions such as: (1) Candidates who graduated before 2015 should be lower on the list, with low priority; (2) Candidates with more years of NLP experience should be ranked higher, with medium priority; and (3) The most over-qualified candidate should appear at the end, with high priority, as the recruiter wants to focus on other candidates first.
Appendix C Details of Prompts
The prompts utilized for ranking token-level and paragraph-level items are detailed in Prompts C and C, respectively. Additionally, the prompts employed for the extraction and sorting of conditions are outlined in Prompts C and C. Finally, we provide an example of zero-shot CoT-based prompt for token-level items in Prompt H.
Item-1: [item-1]
Item-2: [item-1]
Conditions: [string of conditions]
Conditions: [list of extracted conditions]
Appendix D Detailed Breakdown of Ranking performance on MCRank
@@
A detailed breakdowns of models performance on MCRank, segmented by the category of conditions and items, are presented in Tables 8, 9, 10, 11, 12, and 13.
Appendix E LLMs Performances in Satisfying High Priority Conditions
To better understand the relationship between LLMs’ ranking performance and different provided conditions, we calculated the percentage of samples where the “high priority” condition was satisfied. Focusing on high-performing models in MCRank, i.e., OpenAI o1-mini, GPT-4, and Llama3.1, we present the results in Table 5. As shown, the accuracy of satisfying the “high priority” condition changes only slightly as the number of items increases. Additionally, LLMs enhanced with EXSIR show consistently higher accuracy, consistent with their exact match and average ranking accuracy. Notably, the accuracy of satisfying the ’high priority’ condition is significantly higher—especially in GPT-4 and Llama3.1—compared to exact match ranking accuracy, suggesting this condition is only partially responsible for the overall low performance of LLMs on MCRank.
Models Token-level Paragraph-level
3 items 5 items 7 items 3 items 5 items 7 items
o1-mini 34.0 31.1 31.0 38.6 33.1 33.0
o1-mini-Ours 51.6 51.2 47.9 55.2 52.8 49.4
GPT-4 39.4 36.2 35.7 41.3 39.3 36.0
GPT-4-Ours 50.7 50.5 48.3 54.4 47.4 44.9
Llama3.1 37.0 33.6 33.4 43.7 37.4 35.7
Llama3.1-Ours 51.1 49.6 49.1 45.5 43.7 37.1
Table 5: LLMs’ accuracy in satisfying high priority conditions when ranking items in MCRank.
Appendix F Impact of Zero-Shot CoT on Llama3.1
The accuracy of Llama3.1, Llama3.1 with zero-shot CoT, and Llama3.1 with EXSIR on MCRank is shown in Figures 7 and 8. As shown, the figures reveal that while CoT prompting, in most cases, improves Llama3.1’s base performance, there remains a significant performance gap, especially as the number of conditions increases, between EXSIR and CoT, emphasizing the benefits of multi-step reasoning.
Appendix G Impact of Few-Shot CoT on MCRank
While we only present zero-shot CoT results in the paper, we did investigate the performance of few-shot CoT prompting on the MCR task. Our observations showed that few-shot CoT performed only marginally better, and only when the provided demonstrations were fully aligned with the target sample—specifically, when they had the same number and category of conditions, as well as a similar number and type of items (consistent with similar observations reported in (Stechly et al., 2024)). Moreover, since assuming access to demonstrations that perfectly mirror the target sample is unrealistic in real-world scenarios, and to ensure a more consistent evaluation across various cases, we chose to focus primarily on zero-shot CoT results in the paper.
Appendix H More Details on Existing Rankers and o1-mini
To evaluate the performance of encoder-based rankers on MCRank, we treat the instruction string as the query. Then, using LlamaIndex (Liu, 2022), we rank the list of items for each sample individually. The results for ColBERT on MCRank are shown in Figures 7 and 8. We also provided more detailed results of and o1-mini performance in Tables 6 and 7. Moreover, the average accuracy of decomposition step for o1-mini (similar to the results in Table 3) are as follows: for token-level items, 98.4% with 2 conditions and 95.0% with 3 conditions; for paragraph-level items, 96.1% with 2 conditions and 93.2% with 3 conditions.
Figure 7: Evaluating the impact of EXSIR on Llama3.1 against zero-shot CoT prompting for token-level items. We additionally report ColBERT performance.
Figure 8: Evaluating the impact of EXSIR on Llama3.1 against zero-shot CoT prompting for paragraph-level items. We additionally report ColBERT performance.
Models 3 items 5 items 7 items
ACC Avg ACC ACC Avg ACC ACC Avg ACC
Base 1 condition 81.0 81.7 71.4 74.9 66.1 74.5
2 conditions 52.6 56.4 37.1 50.0 23.2 40.1
3 conditions 28.3 38.2 14.4 34.0 8.4 31.8
Ours 1 condition 80.2 81.3 72.1 75.6 65.0 73.8
2 conditions 57.8 61.1 42.3 54.9 26.8 46.2
3 conditions 39.6 47.1 28.8 43.2 17.9 37.6
Table 6: o1-mini and o1-mini with EXSIR (ours) performances for token-level setting in MCRank.
Models 3 items 5 items 7 items
ACC Avg ACC ACC Avg ACC ACC Avg ACC
Base 1 condition 79.5 80.0 74.3 76.1 69.9 72.4
2 conditions 51.4 55.7 33.2 44.9 21.0 38.9
3 conditions 28.9 39.0 14.0 33.5 7.5 30.6
Ours 1 condition 79.0 79.6 73.5 75.4 70.3 72.7
2 conditions 55.7 61.8 41.0 53.8 31.1 49.7
3 conditions 35.4 44.8 17.2 37.9 11.6 33.2
Table 7: o1-mini and o1-mini with EXSIR (ours) performances for paragraph-level setting in MCRank.
@@ Appendix H More Details on Existing Rankers and o1-mini
Llama3.1 Positional 42.0 42.0 47.0 47.0 46.0 46.0
Locational 87.2 87.2 48.3 48.3 83.0 83.0
Temporal 66.5 69.6 50.5 60.3 45.0 55.2
Trait-based 78.0 85.0 40.0 72.7 38.5 77.9
Reason-based 94.0 94.8 94.0 94.0 94.0 94.0
All 73.1 75.4 57.7 69.3 56.6 65.9
@@
Llama3.1-Ours Positional 41.0 41.0 46.1 46.1 46.5 46.5
Locational 85.7 85.7 49.5 49.3 84.4 84.4
Temporal 67.8 71.3 52.2 61.8 43.5 53.4
Trait-based 77.3 84.0 39.2 71.9 36.7 76.1
Reason-based 93.1 93.5 95.0 95.0 92.6 92.6
All 72.6 74.8 58.1 69.7 56.0 65.2
@@
Llama3.1 Positional 39.4 35.4 11.3 30.5 2.7 22.5
Locational 44.3 51.9 12.5 38.0 11.7 34.3
Temporal 38.0 36.1 11.0 23.7 10.5 21.2
Trait-based 29.0 42.6 15.0 42.4 4.5 31.5
Reason-based 42.5 52.9 19.7 41.2 5.6 37.0
All 34.3 43.5 13.8 34.7 6.3 27.7
@@
Llama3.1-Ours Positional 40.1 47.5 25.9 46.3 20.1 38.6
Locational 46.7 53.4 34.8 50.4 13.7 41.8
Temporal 20.0 30.0 11.0 27.8 7.5 22.5
Trait-based 43.0 54.8 19.5 44.5 12.5 41.4
Reason-based 47.0 55.4 22.8 45.4 13.8 47.8
All 39.1 48.1 21.5 41.9 13.2 36.8
@@
Llama3.1 Positional 10.0 31.0 3.6 29.2 0.0 26.2
Locational 25.7 34.8 5.5 31.4 0.0 21.4
Temporal 24.5 37.8 2.5 19.3 0.5 17.4
Trait-based 23.5 38.3 8.5 30.3 4.5 29.5
Reason-based 28.0 38.4 10.3 34.1 2.1 31.4
All 22.3 36.1 5.7 27.9 1.6 24.5
@@
Llama3.1-Ours Positional 5.2 25.5 0.0 20.8 0.0 18.1
Locational 33.3 44.0 19.4 40.1 3.8 28.8
Temporal 30.0 42.9 7.5 31.8 2.0 25.0
Trait-based 44.0 52.3 22.0 44.6 18.0 46.2
Reason-based 36.0 43.5 19.8 36.0 10.6 40.1
All 29.8 41.8 12.9 34.5 7.1 31.3
@@
Llama3.1 Positional 40.5 40.5 44.0 44.0 42.0 42.0
Locational 96.5 96.5 97.0 97.0 95.0 95.0
Temporal 62.0 65.6 46.0 54.0 29.5 40.9
Trait-based 85.0 86.3 79.0 81.1 73.0 74.7
Reason-based 37.5 42.4 19.0 25.4 15.0 20.2
All 64.3 66.3 57.0 60.3 50.9 54.5
@@
Llama3.1-Ours Positional 40.0 40.0 43.0 43.0 42.0 42.0
Locational 97.5 97.5 97.5 97.5 98.0 98.0
Temporal 60.0 63.3 47.0 52.7 32.0 40.4
Trait-based 87.5 88.5 80.5 82.3 70.5 72.2
Reason-based 34.5 40.0 19.5 28.3 20.0 25.4
All 63.9 65.8 57.5 60.7 52.5 55.6
@@
Llama3.1 Positional 25.0 36.2 12.5 30.2 6.0 24.5
Locational 41.0 52.5 25.5 41.9 18.5 36.1
Temporal 26.5 42.2 11.0 30.2 10.0 26.4
Trait-based 34.5 45.8 24.0 39.8 13.0 27.8
Reason-based 24.0 36.4 9.5 26.4 4.0 21.1
All 30.2 42.6 16.5 33.7 10.3 27.2
@@
Llama3.1-Ours Positional 34.0 44.4 16.0 33.0 6.0 27.3
Locational 50.5 60.3 22.5 46.0 21.0 45.0
Temporal 34.5 46.2 11.5 31.0 3.5 19.0
Trait-based 43.0 53.7 26.0 44.1 15.5 29.6
Reason-based 23.0 36.3 8.0 24.9 4.0 16.5
All 37.0 48.2 16.8 35.8 10.0 27.5
@@
Llama3.1 Positional 7.0 27.6 2.0 26.6 1.0 22.0
Locational 33.0 41.9 9.5 35.6 0.0 23.7
Temporal 26.0 42.4 2.5 27.4 0.5 18.5
Trait-based 31.5 42.8 8.0 31.4 0.5 22.4
Reason-based 29.5 43.3 2.5 24.7 0.0 18.7
All 25.4 39.6 4.9 29.1 0.4 21.1
@@
Llama3.1-Ours Positional 29.5 45.8 2.5 27.1 0.0 17.2
Locational 25.5 42.1 10.0 31.5 0.5 18.7
Temporal 25.5 44.4 1.5 24.4 0.0 16.7
Trait-based 26.5 43.1 9.0 36.2 0.0 22.5
Reason-based 22.5 33.5 2.5 23.7 0.5 16.5
All 25.9 41.8 5.1 28.5 0.2 18.3